PART 1) USE A WEIGHTED MOING AVERAGE OF THE following data with weights of 3, 2...

PART 1)

USE A WEIGHTED MOING AVERAGE OF THE following data with weights of 3, 2 (MOST RECENT TO OLDEST).

PART 2)

WHat is the MAPE value for your forecast?

Month Actual DEMAND
1 15
2 10
3 30
4 25
Answer #1

Forecast for Month 3 = (10*3+15*2)/(3+2) = 12

Forecast for Month 4 = (30*3+10*2)/(3+2) = 22

Absolute Forecast Error of Month 3 = |Actual - forecast| = |30-12| = 18

Absolute Forecast Error of Month 4 = |25-22| = 3

Absolute % Error for Month 3 = Absolute Forecast Error/Actual Demand = 18/30 = 0.60 = 60%

Absolute % Error for Month 4 = Absolute Forecast Error/Actual Demand = 3/25 = 0.12 = 12%

MAPE = Average of Absolute % Error = (0.60+0.12)/2 = 0.36 = 36%